problem 3:
problem4:
Problem 3 A manufacturer has four classifications (A, E, LL, and V) for its components. A batch of 100 components is expected to be populated with 45, 25, 4, and 26 of components A, E, LL, and V, respectively. The company is interested in carrying out quality control measures on their products by randomly sampling a certain number of products from each batch. What is the smallest sample size that can be used by the manufacturer to ensure a 40% probability of catching exactly one LL component? (Hint: start with a sample size of at least 10)
Problem 4 While testing the material properties for batches of concrete to be used in building foundations, construction firm estimates that a randomly selected batch will have a 70% chance of passing these a tests on average. What is the probability that the tenth batch tested is the seventh batch to pass? What is the likelihood that the fifth batch is the first one to pass the material property tests?
Answer #1

3)

p = 4/100 = 0.04

we need

1 - (1-p)^n >= 0.4

1 - 0.96^n >= 0.4

n >= 12.51

n = 13 is correct